OpenStudy (zepp):

I need a quick help here, it's about acceleration of a falling object, ignoring air resistance; I have to calculate a baseball's falling speed, just right before it hits the ground, I did everything and found that this ball with start falling from 48.42 meters. (Exactly at \(\frac{4745}{98}\) So from what I understood from class and book, this object has no more speed when it starts to fall, therefore, no more kinetic energy and has maximum potential energy. I thought of find the speed with the formula \(\large E_k = \frac{1}{2}mv^2\) but I'm really not sure about it;

OpenStudy (zepp):

So I tried to do it with the Earth's acceleration, which is \(\large 9.8\frac{m}{s^2}\)

OpenStudy (zepp):

but I'm don't know how, can someone guide me through this please? Thanks :)

OpenStudy (zepp):

@KingGeorge @cshalvey Can you help me please? :D

OpenStudy (kinggeorge):

Have you tried using \(-9.8 m/s^2\)?

OpenStudy (zepp):

Nope, I don't know how it works D:

OpenStudy (kinggeorge):

Well, if you're asked for speed, you should use the absolute value of velocity, and your way appears to be valid. However, if you're asked about velocity, you should use \[-9.8\frac{m}{s^2}\]as your acceleration instead.

OpenStudy (kinggeorge):

Let me see what I get when I use a negative acceleration.

OpenStudy (kinggeorge):

I'm getting a speed of \(\approx30.8064\) and a velocity of about -30.8064

OpenStudy (zepp):

I got that as well, with \(\large E_k = \frac{1}{2}mv^2\)

OpenStudy (zepp):

but can you show me how to got 30.8064 with the acceleration factor? :D

OpenStudy (kinggeorge):

Use the formula \[v_f^2=v_0^2+2ad\]Where \(v_0\) is the initial velocity, \(a\) is acceleration, and \(d\) is the distance traveled.

OpenStudy (kinggeorge):

I've got to go to dinner now, but I'll be back to answer more questions later

OpenStudy (zepp):

Alright, I think I've understood that acceleration thing now, thank you very much! :]

OpenStudy (stormfire1):

An easy way to calculate this is to simply use:\[PE_{top}=mgh\]\[KE_{ground}=\frac{1}{2}mv^2\]Conservation of energy states that PE will equal KE in this situation so:\[mgh=\frac{1}{2}mv^2\]Simplifying you get: \[\not mgh=\frac{1}{2}\not mv^2\]\[(9.8m/s^2)(48.42m)=\frac{1}{2}v^2\]Now just solve for v:\[v=\sqrt{(2)(9.8m/s^2)(48.42m)}\]\[v=30.8m/s\]

OpenStudy (stormfire1):

Note at the top that I meant \[\Delta KE~will~equal~ \Delta PE\]

OpenStudy (zepp):

Never thought of conservation. It's all clear now, Thanks a lot! :D

OpenStudy (stormfire1):

Cool. Also note that if you consider downward negative in this problem the final value will obviously be -30.8 m/s.

OpenStudy (zepp):

Yes, I got that part, thanks to KingGeorge, thank you both! :)